# Harnessing Knowledge Retrieval with Large Language Models for Clinical Report Error Correction

### v1

@@ Appendix A Appendix
Experiment Setups.
All the experiments are conducted with 4 NVIDIA GeForce RTX 3090. For our LLM experiments, we utilize the AutoModelForCausalLM from the Hugging Face Transformers library. The model is loaded from the specified model_id, with the torch_dtype parameter set to torch.bfloat16. We set the max_new_tokens=300, do_sample=True, temperature=0.001 and top_p=0.8, all other hyperparameters remains unchanged as default values.
RAG pipeline.
Evaluation Data Preparation.
The list of 12 observations we choose to replace: ’Atelectasis’, ’Cardiomegaly’, ’Consolidation’, ’Edema’, ’Enlarged Cardiomediastinum’, ’Fracture’, ’Lung Lesion’, ’Lung Opacity’, ’Pleural Effusion’, ’Pleural Other’, ’Pneumonia’, ’Pneumothorax’.
The list of replacement words are: ’Asthma’,’Costochondritis’,’Pulmonary Embolism’, ’Thoracic Outlet Syndrome’,’Tracheitis’,’Tracheomalacia’,’Vocal Cord Dysfunction’, ’Pharyngitis’,’Laryngitis’,’Mesothelioma’,’Obstructive Sleep Apnea’, ’Aspergillosis’; ’Appendicitis’, ’Gastroesophageal Reflux Disease’, ’Crohns Disease’, ’Ulcerative Colitis’, ’Gallstones’, ’Pancreatitis’, ’Hepatitis’, ’Cirrhosis’, ’Irritable Bowel Syndrome’, ’Peptic Ulcer’, ’Celiac Disease’, ’Diverticulitis’, ’Hemorrhoids’, ’Anal Fissure’, ’Intestinal Obstruction’, ’Inflammatory Bowel Disease’, ’Gastroparesis’, ’Cholecystitis’, ’Gastric Ulcer’, ’Duodenal Ulcer’, ’Esophageal Varices’, ’Achalasia’, ’Barretts Esophagus’, ’Esophageal Cancer’, ’Pancreatic Cancer’, ’Colorectal Cancer’, ’Liver Cancer’, ’Gastric Cancer’, ’Hiatal Hernia’, ’Esophageal Stricture’.

### v2

@@ Appendix A Appendix
Experiment Setups.
All the experiments are conducted with 4 NVIDIA GeForce RTX 3090 24576MiB. For our LLM experiments, we utilize the AutoModelForCausalLM from the Hugging Face Transformers library. The model is loaded from the specified model_id, with the torch_dtype parameter set to torch.bfloat16. We set the max_new_tokens=300, do_sample=True, temperature=0.001 and top_p=0.8, all other hyperparameters remains unchanged as default values.
RAG Pipeline.
Evaluation Data Preparation.
Table 8 and 9 shows the words used for extracting and replacing in the original reports to construct the benchmark data. We got the list from radiologists where we chose 12 important observation entities to replace, and 41 words that is not relevant to the report as the replacement words.
Words to Replace
Atelectasis
Cardiomegaly
Consolidation
Edema
Enlarged Cardiomediastinum
Fracture
Lung Lesion
Lung Opacity
Pleural Effusion
Pleural Other
Pneumonia
Pneumothorax
Table 8: Words to replace
Replacement Words
Asthma; Costochondritis; Pulmonary Embolism;
Thoracic Outlet Syndrome; Tracheitis;
Tracheomalacia; Vocal Cord Dysfunction;
Pharyngitis; Laryngitis; Mesothelioma;
Obstructive Sleep Apnea; Aspergillosis;
Appendicitis; Gastroesophageal Reflux Disease;
Crohn’s Disease; Ulcerative Colitis; Gallstones;
Pancreatitis; Hepatitis; Cirrhosis; Peptic Ulcer;
Irritable Bowel Syndrome; Celiac Disease
Diverticulitis; Hemorrhoids; Anal Fissure
Intestinal Obstruction; Gastroparesis
Cholecystitis; Gastric Ulcer; Duodenal Ulcer
Esophageal Varices; Achalasia; Barrett’s Esophagus
Esophageal Cancer; Pancreatic Cancer;
Inflammatory Bowel Disease; Colorectal Cancer
Liver Cancer; Gastric Cancer; Hiatal Hernia;
Esophageal Stricture
Table 9: Replacement words.
@@
Human Evaluation
Applying LLMs in the medical domain may not be entirely convincing if relying solely on automated evaluation metrics. This is because automatic evaluation metrics primarily focus on assessing semantic similarity and sentence structure, while in the medical field, even minor errors in generated text can have severe consequences. To provide a more comprehensive evaluation of the effectiveness of our methods, we conducted a human assessment one the generated tasks, which is the error correction task. We choose 50 examples with six models across the three different inference approaches for comparison: end-to-end, three-stage inference (baseline), and three-stage inference (Internal + Ex). This allowed us to evaluate the performance of the models in a more nuanced manner beyond relying solely on automated evaluation metrics. Detailed description of the evaluation criteria can be found in Table 11. We use a Likert scale ranging from 1 to 5 for the evaluation
Metrics
Accuracy: The model correctly identifies and addresses errors in the medical report.
- Does the model accurately identify all errors in the report?
- Are the corrections made by the model accurate and appropriate?
Factual Consistency: The model’s output maintains consistency with established medical facts and does not introduce incorrect information.
- Does the corrected information align with current medical knowledge?
- Are there any instances of contradictory or inconsistent information in the corrections?
Clinical Relevance: The model’s corrections and additions are clinically appropriate and enhance the overall quality and usefulness of the medical report.
- Do the corrections improve the clinical value of the report?
- Are the changes relevant to the patient’s condition and potential treatment?
Table 11: Metrics used for human evaluation.
Figure 3 presents the human evaluation scores for six models across three critical dimensions: Accuracy, Factual Consistency, and Clinical Relevance. The results reveal notable performance differences among the evaluated methods.
The three-stage inference (baseline) method consistently outperforms the end-to-end approach, albeit by a small margin. This slight edge is evident across all three dimensions for most models. During the human evaluation process, we observed that outputs from models using the three-stage inference (baseline) method exhibited more complete and coherent report structures. This observation may explain why this method’s automatic evaluation metrics were higher than those of the end-to-end approach, despite the relatively small difference in human evaluation scores.
Most strikingly, the three-stage inference (internal + external) method demonstrates significant improvements in model performance across all three dimensions. This method consistently outperforms other inference approaches, with all models scoring notably higher in accuracy, factual consistency, and clinical relevance. The improvement is particularly pronounced for models like Phi3-medium and Llama3-8B, where the gains are substantial across all metrics. Interestingly, the magnitude of improvement varies across models. For instance, the MMedLM2 model shows a dramatic increase in performance when using the three-stage inference (internal + external) method, especially in terms of factual consistency and clinical relevance. In contrast, the Aloe model exhibits a more modest, though still significant, improvement.
The reason why the automatic evaluation results did not show a highly significant improvement for the three-stage inference (internal + external) method compared to the three-stage inference (baseline) method may be that automatic evaluation metrics primarily focus on structure and overall semantic similarity. These metrics may not effectively capture the correct identification and correction of specific errors in the reports, which are crucial aspects that human evaluators can discern.
This discrepancy between automatic and human evaluation results underscores the importance of human assessment in evaluating the quality of medical report generation models. It suggests that while automatic metrics are useful for initial screening and large-scale evaluations, they may not fully capture the nuanced improvements in accuracy, factual consistency, and clinical relevance that are critical in medical contexts.
Case Study
Figure 2: Case study. Our method refers to the three-stage inference with internal + external retrieval. We highlight the correct part in red and the error part in blue.
The case study (Figure 2) demonstrates significant improvements in error detection and correction capabilities using our method across various models. Among the six models examined, Llama3-8b, Phi3-small, and Phi3-medium successfully corrected the error, showcasing the effectiveness of our approach. Notably, Phi3-medium’s correction perfectly matched the ground truth, while Llama3-8b and Phi3-small provided semantically correct corrections with slight variations in expression. Compared to the end-to-end approach, our method enabled Aloe and Phi3-mini to successfully identify the erroneous sentences, although they fell short in providing accurate corrections. This improvement in error detection, even without perfect correction, highlights the enhanced capabilities our method brings to these models.
Figure 3: Human evaluation comparison.